OpenStudy (anonymous):

a box has dimension of 10-x, 10-x and 1/2x. Write an expression in terms of x for the volume of the box. b. use the expression to calculate the volume of x.. c. what is the maximum volume of the box?

OpenStudy (pfenn1):

The volume(V) of the box would be L*W*H where L is the length 10-x W is the width 10-x H is the height 1/2 x V=(x/2)(10-x)^2

OpenStudy (pfenn1):

I don't understand what "calculate the volume of x" means?

OpenStudy (anonymous):

x must be the vollume of the box. How would you calculate the maximum volume?

OpenStudy (pfenn1):

I am not sure. I would assume that x should be small so that 10-x would be as big as possible but I am not really sure.

OpenStudy (anonymous):

ok. so say ur assumption is correct. what would be the final result

OpenStudy (pfenn1):

do you think the height is (1/2)x or 1/(2x)? If it is the first, then the volume will go down as x gets smaller. in the second. the volume goes up as x gets smaller

OpenStudy (anonymous):

i think its 1/(2x)

OpenStudy (pfenn1):

The maximum volume would be when x becomes infinitely small, that would make the volume infinitely big. hmm. you might need to find someone who gets this better than me.

Directrix (directrix):

If the volume of the rectangular solid container is V(x) = ((x/2)(10-x)^2), then the maximum volume of the box would occur at x = 10/3 yielding a maximum volume of V(10/3) = 2000/2. http://www.wolframalpha.com/input/?i=y+%3D%28x%2F2%29%2810-x%29%5E2